Write an equation in slope-intercept form for the line that is perpendicular to the given line and that passes through the given point. 3x - 4y = 8; (-6,5)
Answer by robertb(5830)

The equation should be of the form 4x + 3y = c, where c is an unknown constant. To find it, use the coordinates of the point given:
4*-6 + 3*5 = -24 + 15 = -9. Therefore the equation of the line is 4x + 3y = -9. In slope-intercept form, 3y = -4x-9, or y = -4x/3 - 3.